# Stationary densities and delocalized domain walls in asymmetric exclusion processes competing for finite pools of resources

### v1

@@ Abstract
We explore the stationary densities and domain walls in the steady states of a pair of asymmetric exclusion processes (TASEPs) antiparallelly coupled to two particle reservoirs without any spatial extent by using the model in Haldar et al, Phys. Rev. E 111, 014154 (2025). We show that the model admits delocalized domain walls which exist for some choices of the model parameters that define the effective entry and exit rates into the TASEP lanes. Surprisingly, in the parameter space spanned by these model parameters, the delocalized domain walls covers an extended region. The corresponding phase diagrams in the plane of the control parameters have different topology from those for an open TASEP or other models with multiple TASEPs connected to two reservoirs.
@@ I INTRODUCTION
Driven diffusive systems driv1 ; driv2 ; driv3 ; driv4 can show highly nontrivial properties even with purely local dynamics in even one dimension (1D). A paradigmatic example is the totally asymmetric simple exclusion process (TASEP) krug-prl ; derrida1 ; derrida2 ; derrida3 . It consists of a 1D lattice with open boundaries. A TASEP’s dynamics are stochastic and subject to exclusion at all sites, meaning that it can only hold a single particle per site. A particle is entered at one of the boundaries at a given rate, hops unidirectionally to the subsequent sites at a rate of unity until it reaches the other end, and then leaves the system at another specified rate. The TASEP shows three different phases as a function of the entry and exit rates at the boundary: the low-density phase is given by steady-state densities in bulk that are less than half, the high-density phase by densities greater than half, or the maximal current phase with densities of just half.
@@
The TASEP was first put forth as a minimal model to explain how proteins are synthesized in biological cells macdonald . Subsequently it was found to display boundary-induced phase transitions krug-prl ; derrida1 ; derrida2 . TASEP has been generalized to model various phenomena in biological and social systems, which has resulted in the identification of new phases and novel collective phenomena. For example, it has been found that traffic jams can be caused by the particle exchange between the environment and the TASEP lane ef-lktasep-prl ; ef-lktasep-pre , which has been confirmed experimentally traffic-exp . Further investigation of these findings revealed that molecular motors belonging to the kinesin-8 family can govern the kinetics of microtubule depolymerization, which in turn leads to microtubule length regulation melbinger1 ; melbinger2 . Furthermore, systems with two TASEP lanes coupled by infrequent particle-switching events have been found to exhibit complex phenomena, e.g., domain wall delocalization and traffic jams reichenbach1 ; reichenbach2 ; reichenbach3 .
@@
TASEP has also been investigated on simple networks, examining the requirements for the existence of the different phases and domain walls net1 ; net2 ; net3 on different segments of the networks. These studies have been inspired by the dynamics of molecular motors traversing intricate networks of microtubules in real cells net-exp . In yet another generalization of TASEP, the interaction of particle number conservation with TASEP on an inhomogeneous ring has been explored and found to produce macroscopically nonuniform stable states of various kinds def1 ; def2 ; sarkar-basu ; banerjee-sarkar-basu ; def5 ; def6 ; def7 ; def8 ; def9 ; def10 . Combining TASEP with the symmetric exclusion process (SEP) provides yet another intriguing generalization of TASEP. Both half-closed half1 ; half2 and fully open systems half3 have been used in this context. These studies have produced some noteworthy findings, including complex space-dependent stationary states in both TASEP and SEP lanes half3 , self-organized temporal patterns half2 , and tip localization of motors half1 . In approaches complementary to Refs. ef-lktasep-prl ; ef-lktasep-pre , a TASEP lane is considered to be connected to an embedded three-dimensional (3D) particle reservoir via particle exchanges in order to mimic the coupled system of molecular motors and microtubules. For instance, Refs. klumpp1 ; klumpp2 ; ciandrini ; dauloudet have examined the effects of diffusion of the motors on the filament’s steady states (TASEP), providing physical insights into the nontrivial steady states that result from combining 1D driven (TASEP) and 3D equilibrium (diffusion) processes. In addition to these investigations, but adhering strictly to a 1D description, Ref. klumpp3 examined a 1D open model with serially coupled diffusive and driven segments and the alterations that resulted in the pure open TASEP phase diagram. TASEP has also been used as a minimal model for traffic flow along network of roads Chowdhury-Santen-Schadschneider ; Schadschneider .
@@
In this work, we study a distinct class of TASEP models that are very different from those that have been discussed above. These models place a limit on the total number of particles in the combined system by connecting the TASEP to a particle reservoir at both ends rather than having open boundaries. These models are inspired by the fact that in the real world, limited amount of resources are available for use. For instance, finite availability of ribosomes in mRNA translocation for protein synthesis in eukaryotic cells adams-schmittmann-zia , or a finite number of vehicles in a closed network ha-den . This necessitates the extension of an unconstrained TASEP into systems that are composed of TASEP(s) connected to particle reservoir(s) with overall particle conserving dynamics. These models can also describe traffic dynamics, where a network of roads having traffic flow contains a limited fixed number of vehicles traff1 ; see Ref. traff2 for similar studies. In these models, the primary effect of the finite availability of resources is that the effective entry rate of the particles to the TASEP, e.g., the actual protein synthesis taking place, sensitively depends upon the available resources. As the resources go up (reduce), protein synthesis rates increase (decrease). The main qualitative question in each of these studies is how the steady-state occupations and currents differ from open TASEPs due to the limited pool of accessible particles. From the standpoint of nonequilibrium statistical mechanics, these models serve as minimal models for “boundary-induced nonequilibrium phase transitions with a conservation law”. In the recent past, there have been several studies focusing on aspects of the stationary states in TASEPs with finite resources systems and are composed of single TASEP or more ha-den ; adams-schmittmann-zia ; ciandrini ; cook-dong-lafleur connected to a particle reservoir. How the phase diagrams and stationary densities and domain walls are controlled by the finite resources is the major theme of these works. In all these previous studies, the effective entry rate depends on the instantaneous reservoir population, however the particle exit rates from the TASEP lane to the reservoir are considered a constant, independent of reservoir population. Moreover, only one reservoir is employed in all these studies. Considering reservoir population-dependent exit rates allow one to study the “reservoir crowding effect”, proposed and studied in Ref. astik-parna . The role and effects of distributed resources, as opposed to a single point reservoir, has also been explored in the recent past. These studies include research on how two reservoirs interact with two TASEPs sourav-1 ; sourav-2 ; astik-erwin and how reservoirs interact in a bidirectional system arvind-bi . See Ref. seppa for a related study on a TASEP with $K$-exclusion, which generalizes the conventional TASEP and allows for up to $K$ particles at each site.
@@
In this article, we explore the stationary densities and the domain walls in a model with two asymmetric exclusion processes coupled antiparallelly to two particle reservoirs. This was originally proposed and studied in Ref. astik-erwin . By considering a reduced parameter space distinct from that considered in Ref. astik-erwin , we show that the stationary densities in the two TASEP lanes are statistically identical, which can be spatially uniform or nonuniform. In the latter case, we find a pair of delocalized domain walls (DDW), one in each of the TASEP lanes, spanning completely or partially the lanes. The phase diagrams of the system in this reduced parameter space as functions of the control parameters are obtained, which are quite different from those found in Ref. astik-erwin , or for other models of TASEPs with finite resources. The most striking feature is that the phase space region in the plane of the control parameters where the DDWs are found is extended, as opposed to just a line found usually in other models of TASEP with or without open boundaries, or with finite resources; see, e.g.,  blythe ; sarkar-basu ; astik-parna ; astik-erwin among others. Apart from its theoretical novelty, this has profound phenomenological implications: it means that for the control parameters in an extended region of the phase space, the stationary densities actually have large fluctuations, since DDWs are nothing but moving density shocks. The remainder of this article is organized as follows: In Section II, we discuss the model that we study. Then in Section III, we outline the construction of the MFT for this model. Next in Section IV, we present the phase diagrams in the plane of the control parameters. Then in Section V, we use MFT to obtain the stationary densities in the different allowed phases and corroborate them by extensive MCS studies, and also discuss the absence of nonidentical stationary densities in the two TASEP lanes. Then in Sec. VI, we calculate the different phase boundaries, and the nature of the phase transitions in the model. We conclude by providing a short summary of our results and future prospects in Section VII. Additional results and some technical details are available in Appendix A and Appendix B for interested readers.
@@ II THE MODEL
We briefly revisit the construction and dynamics of the model proposed and studied in Ref. astik-erwin that we use for our work. The model is composed of two 1D lattices, $T_{1}$ and $T_{2}$, of equal length exhibiting dynamical rules of TASEP which are connected to two reservoirs, $R_{1}$ and $R_{2}$, with finite particle-carrying capacity at their ends, see Fig. 1 for the schematic diagram. The TASEP lanes are of equal length $L$ with thier sites labeled as j=1,2,,Lj=1,2,...,L. Assumed to be point reservoirs without any spatial extent or internal dynamics, $R_{1}$ and $R_{2}$ respectively have $N_{1}$ and $N_{2}$ number of particles in them that are functions of time. The particles in the system interact with each other only through hard-core exclusion, meaning that each site on both TASEP lanes can be occupied by no more than one particle at a given time. Thus, each TASEP site has only two states: either empty or occupied. We follow the dynamics prescribed in Ref. astik-erwin . We briefly describe that below for the convenience of the reader.
@@
    The effective rates depend on the instantaneous reservoir populations $N_{1}$ and $N_{2}$ (Eqs. (2)-(4)).
@@
As in Ref. astik-erwin , the effective entry and exit rates of particles into and out of the TASEP lanes are parametrized by the parameters $\alpha_{i}$ and $\beta_{i}$ ($i=1,2$), which are positive definite, as follows:
@@
The functions $f(N_{i})$ and $g(N_{i})$ must be monotonically increasing and decreasing in $N_{i}$, respectively, since an increasing reservoir population promotes particle entry into the associated TASEP lane but inhibits their exit. Various forms of such rate functions satisfying monotonicity have been proposed in the literature greulich-ciandrini ; adams-schmittmann-zia ; cook-dong-lafleur ; astik-parna ; arvind1 ; arvind2 ; sourav-1 ; sourav-2 . For reasons of simplicity, we use the forms in Ref. astik-erwin
@@
To reduce the number of control parameters, we consider the symmetric case $\alpha_{1}=\alpha_{2}\equiv\alpha$ and $\beta_{1}=\beta_{2}\equiv\beta$. These choices are distinct from those in Ref. astik-erwin , and allow us to explore a different reduced parameter space from the one considered in Ref. astik-erwin . The model is thus governed by three parameters: $\alpha$, $\beta$, and $\mu$.
@@ III MEAN-FIELD THEORY
In this section, we analyze the model at the mean-field blythe level, following the logic developed in Ref. astik-erwin .
@@
Eqs. (6), (7), and (8) are invariant under the following set of transformations (see Appendix A for details):
@@
To move toward a continuum description, we introduce a coarse-grained spatial variable $x=j\epsilon$, with lattice spacing $\epsilon=1/L$. In the limit $L\gg 1$, $x$ becomes quasi-continuous, and the density $\rho^{(i)}(x,t)$ is defined accordingly.
@@
becomes constant. Denoting the constant currents in lanes $T_{1}$ and $T_{2}$ as $J^{(1)}$ and $J^{(2)}$ respectively, the flux-balance-equation relating these currents to the reservoir occupation numbers $N_{1}$ and $N_{2}$ reads:
@@ IV Mean-field phase diagrams of the model for $0<\mu\leq 1$
Before we analyze the phase diagrams of the present model, it is useful to review of the phase diagram of an open TASEP schmittmann-zia ; chou-mallick-zia ; blythe . In an open TASEP, depending on the entry or exit rates ($\alpha_{T}$ and $\beta_{T}$), one finds LD, HD, and MC phases with (constant) steady state densities $\rho_{\text{LD}}=\alpha_{T}$, $\rho_{\text{HD}}=1-\beta_{T}$, and $\rho_{\text{MC}}=1/2$, respectively, in the bulk. Conditions for obtaining these phases are: $\alpha_{T}<\beta_{T}$ and $\alpha_{T}<1/2$ (LD phase), $\beta_{T}<\alpha_{T}$ and $\beta_{T}<1/2$ (HD phase), and $\alpha_{T},\beta_{T}\geq 1/2$ (MC phase). In addition to these, a delocalized domain wall (DDW) is also formed for $\alpha_{T}=\beta_{T}<1/2$. This DDW is manifested as an inclined line spanning the entire TASEP lane and is a consequence of the particle nonconserving dynamics in an open TASEP. Now, transitions between different phases can be identified by equating the particle currents in the associated phases. Thus, the boundaries between either LD or HD and MC phases come out to be straight lines parallel to the $\beta_{T}$ and $\alpha_{T}$ axes respectively, wherein the LD and HD phases are separated by a coexistence line (CL) which starts at (0,0) and terminates at (1/2,1/2) in the $\alpha_{T}$-$\beta_{T}$ plane. Considering the bulk density, $\rho$, as the order parameter, one finds the transition from LD to HD phase as discontinuous change in $\rho$, thereby implying it a discontinuous phase transition. Contrary to this, the transition from either LD or HD to MC phase is seen to be a continuous change in $\rho$, implying it a continuous phase transition.
@@
We now illustrate the phase diagrams of our model in the plane of control parameters $\alpha$ and $\beta$ for a set of representative values of filling factor $\mu$ (cf. Fig. 2 and 8). We use mean-field theory to obtain the phase diagrams and density profiles, which are corroborated by an extensive Monte Carlo simulation (MCS) studies with random updates. MFT and MCS results agree well with each other. In our model, depending on the value of $\mu$, two or four distinct phases appear simultaneously. These are LD-LD, HD-HD, MC-MC and DW-DW phases. Interestingly, the domain wall solutions, which form the DW-DW phase in our model, always emerge in pair for a particular values of $\alpha$, $\beta$, and $\mu$ are actually moving density shocks, not pinned to specific locations in the TASEP lanes, having envelops span a finite region across the lanes. These are the delocalized domain walls (DDW) mentioned above. For small values of $\mu$, such as $\mu=3/8$ (Fig. 2(a)), only the LD-LD and DW-DW phases are present, separated by a curved phase boundary. This structure persists for all $0<\mu<1/2$, with the LD-LD region gradually shrinking as $\mu$ increases. In this regime, the HD-HD and MC-MC phases are absent. As $\mu$ increases beyond $1/2$, the MC-MC and HD-HD phases start to emerge, for instance at $\mu=5/8$ (Fig. 2(b)). The LD-LD and HD-HD regions now share boundaries with the MC-MC phase, and these boundaries are straight lines parallel to the $\beta$ and $\alpha$ axes, respectively, independent of the value of $\mu$. This topological structure remains intact at $\mu=7/8$ (Fig. 2(c)), where the HD-HD region continues to grow while the LD-LD and DW-DW regions contract. At the half-filling point ($\mu=1$) (Fig. 2(d)), the phase diagram simplifies further. All the phase boundaries become straight lines, aligned with the coordinate axes in the $\alpha$-$\beta$ phase diagram. Moreover, the particle-hole symmetry of the model is used to construct the phase diagrams above half-filling ($\mu>1$) from their counterparts below half-filling; see Fig. 8 in Appendix A. From these, it is evident that as $\mu$ increases, the LD-LD region diminishes while the HD-HD and DW-DW regions expand. For $\mu>3/2$, both the LD-LD and MC-MC phases vanish, leaving only the HD-HD and DW-DW phases, a trend that continues up to $\mu=2$. Thus, for $1/2<\mu<3/2$, all four phases—LD-LD, HD-HD, DW-DW, and MC-MC—coexist and meet at a single point known as the four-phase multicritical point.

### v2

@@ Abstract
We explore the stationary densities and domain walls in the steady states of a pair of asymmetric exclusion processes (TASEP) antiparallelly coupled to two particle reservoirs without any spatial extent by using the model in Haldar et al., Phys. Rev. E 111, 014154 (2025). We show that the model admits a pair of delocalized domain walls, which exist for some choices of the model parameters that define the effective entry and exit rates into the TASEP lanes. Surprisingly, in the parameter space spanned by these model parameters, the region corresponding to delocalized domain walls covers an extended region, in contrast to the delocalized domain walls that appear only along a line in the relevant parameter space of the other known variants of TASEP. This implies large fluctuations in the TASEP particle numbers even in the thermodynamic limit that can be found over a range of the control parameters. The corresponding phase diagrams in the plane of the control parameters have different topology from those for an open TASEP or other models with multiple TASEPs connected to two reservoirs.
@@ I INTRODUCTION
Driven diffusive systems [1, 2, 3, 4] can show highly nontrivial properties even with purely local dynamics in even one dimension (1D). A paradigmatic example is the totally asymmetric simple exclusion process (TASEP) [5, 6, 7, 8]. It consists of a 1D lattice with open boundaries. A TASEP’s dynamics are stochastic and subject to exclusion at all sites, meaning that it can only hold a single particle per site at a time. A particle enters at one of the boundaries at a given rate, hops unidirectionally to the subsequent sites at a rate of unity until it reaches the other end, and then leaves the system at another specified rate. The TASEP shows three different phases as a function of the entry and exit rates at the boundary: the low-density phase is given by steady-state densities in bulk that are less than half, the high-density phase by densities greater than half, and the maximal current phase with densities of just half.
@@
The TASEP was first put forth as a minimal model to describe proteins synthesis in biological cells [9]. Subsequently it was found to display boundary-induced phase transitions [5, 6, 7]. TASEP has been generalized to model various phenomena in biological and social systems, which has resulted in the identification of new phases and novel collective phenomena. For example, it has been found that traffic jams can be caused by the particle exchange between the environment and the TASEP lane [10, 11], which has been confirmed experimentally [12]. Further investigation of these findings revealed that molecular motors belonging to the kinesin-8 family can govern the kinetics of microtubule depolymerization, which in turn leads to microtubule length regulation [13, 14]. Furthermore, systems with two TASEP lanes coupled by infrequent particle-switching events have been found to exhibit complex phenomena, e.g., domain wall delocalization and traffic jams [15, 16, 17].
@@
TASEP has also been investigated on simple networks, examining the requirements for the existence of the different phases and domain walls [18, 19, 20] on different segments of the networks. These studies have been inspired by the dynamics of molecular motors traversing intricate networks of microtubules in real cells [21]. In yet another generalization of TASEP, the interaction of particle number conservation with TASEP on an inhomogeneous ring has been explored and found to produce macroscopically nonuniform stable states of various kinds [22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Combining TASEP with the symmetric exclusion process (SEP) provides yet another intriguing generalization of TASEP. Both half-closed [32, 33] and fully open systems [34] have been used in this context. These studies have produced some noteworthy findings, including complex space-dependent stationary states in both TASEP and SEP lanes [34], self-organized temporal patterns [33], and tip localization of motors [32]. In approaches complementary to Refs. [10, 11], a TASEP lane is considered to be connected to an embedded three-dimensional (3D) particle reservoir via particle exchanges in order to mimic the coupled system of molecular motors and microtubules. For instance, Refs. [35, 36, 37, 38] have examined the effects of diffusion of the motors on the filament’s steady states (TASEP), providing physical insights into the nontrivial steady states that result from combining 1D driven (TASEP) and 3D equilibrium (diffusion) processes. In addition to these investigations, but adhering strictly to a 1D description, Ref. [39] examined a 1D open model with serially coupled diffusive and driven segments and the alterations that resulted in the pure open TASEP phase diagram. TASEP has also been used as a minimal model for traffic flow along network of roads [40, 41].
@@
In this work, we study a distinct class of TASEP models that are very different from those discussed above. These models place a limit on the total number of particles in the system of the TASEP connected to a particle reservoir at both ends rather than having open boundaries. These models are inspired by the fact that in the real world, limited amount of resources are available for use. For instance, finite availability of ribosomes in mRNA translocation for protein synthesis in eukaryotic cells [42], or a finite number of vehicles in a closed network [43]. This necessitates the extension of an unconstrained TASEP into closed systems that are composed of TASEP(s) connected to particle reservoir(s) with overall particle conserving dynamics. These models can also describe traffic dynamics, where a network of roads having traffic flow contains a limited fixed number of vehicles [44]; see Ref. [45] for similar studies. In these models, the primary effect of the finite availability of resources is that the effective entry rate of the particles to the TASEP, e.g., the actual protein synthesis taking place, sensitively depends upon the available resources. As the resource supply go up (reduce), protein synthesis rates increase (decrease). The main qualitative question in each of these studies is how the presence of a limited pool of accessible particles alters the steady-state densities and currents compared to those of open TASEPs with an unconstrained particle supply. From the standpoint of nonequilibrium statistical mechanics, these models serve as minimal models for “boundary-induced nonequilibrium phase transitions with a conservation law”. In the recent past, there have been several studies focusing on aspects of the stationary states in TASEPs with finite resources systems and are composed of single TASEP or more [43, 42, 37, 46, 47, 48] connected to a particle reservoir. How the phase diagrams and stationary densities and domain walls are controlled by the finite resources is the major theme of these works. In all these previous studies, the effective entry rate depends on the instantaneous reservoir population, however the particle exit rates from the TASEP lane to the reservoir are considered a constant, independent of reservoir population. Moreover, only one reservoir is employed in all these studies. Considering reservoir population-dependent exit rates allow one to study the “reservoir crowding effect”, proposed and studied in Ref. [49]. The role and effects of distributed resources, as opposed to a single point reservoir, has also been explored in the recent past. These studies include research on how two reservoirs interact with two TASEPs [50, 51, 52, 53] and how reservoirs interact in a bidirectional system [54]. See Ref. [55] for a related study on a TASEP with $K$-exclusion, which generalizes the conventional TASEP and allows for up to $K$ particles at each site.
@@
In this article, we explore the stationary densities and the domain walls in a model with two asymmetric exclusion processes (TASEP) coupled antiparallelly to two particle reservoirs. This model was originally considered in Ref. [53]. By considering a reduced parameter space distinct from that considered in Ref. [53], we show that the stationary densities in the two TASEP lanes are statistically identical, which can be spatially uniform or nonuniform. In the latter case, we find a pair of delocalized domain walls (DDW), one in each of the TASEP lanes, spanning completely or partially the lanes. The phase diagrams of the system in this reduced parameter space as functions of the control parameters are obtained, which are quite different from those found in Ref. [53], or for other models of TASEPs with finite resources. The most striking feature is that the phase space region in the plane of the control parameters where the DDWs are found is extended, as opposed to just a line found usually in other models of TASEP with or without open boundaries, or even with finite resources; see, e.g.,  [24, 49, 53, 56] among others. Apart from its theoretical novelty, this has profound phenomenological implications: it means that for the control parameters in an extended region of the phase space, the stationary densities actually have large fluctuations, since DDWs are nothing but moving density shocks. Surprisingly, there is never population imbalances between the two reservoirs, unlike in Ref. [53]. Furthermore, even when the TASEP lanes display large fluctuations in the thermodynamic limit, the relative fluctuations in the reservoir populations decreases monotonically with larger TASEP size, suggesting vanishing relative reservoir number fluctuations in the thermodynamic limit. The remainder of this article is organized as follows: In Section II, we discuss the model. Then in Section III, we outline the construction of the MFT for this model. Next in Section IV, we present the phase diagrams in the plane of the control parameters. Then in Section V, we use MFT to obtain the stationary densities in the different phases and corroborate them by extensive MCS studies. Next, we discuss the absence of nonidentical stationary densities in the two TASEP lanes in Sec. V.5. Then in Sec. VI, we calculate the different phase boundaries, and the nature of the phase transitions in the model. We conclude by providing a short summary of our results and future prospects in Section VII. Additional results and some technical details are available in Appendix A and Appendix B for interested readers.
@@ II THE MODEL
We now revisit the construction and dynamics of the model proposed and studied in Ref. [53] that we use for our work. The model is composed of two 1D lattices, $T_{1}$ and $T_{2}$, of equal length exhibiting dynamical rules of TASEP which are connected to two reservoirs, $R_{1}$ and $R_{2}$, with finite particle-carrying capacity at their ends, see Fig. 1 for the schematic diagram. The TASEP lanes are of equal length $L$ with their sites labeled as j=1,2,,Lj=1,2,...,L. Assumed to be point reservoirs without any spatial extent or internal dynamics, $R_{1}$ and $R_{2}$ respectively have $N_{1}$ and $N_{2}$ number of particles in them that are functions of time. The particles in the system interact with each other only through hard-core exclusion, meaning that each site on both TASEP lanes can be occupied by no more than one particle at a given time. Thus, each TASEP site can be in either of two states: empty or occupied. We follow the dynamics prescribed in Ref. [53]. We briefly describe that below for the convenience of the reader.
@@
    The effective rates depend on the instantaneous reservoir populations $N_{1}$ and $N_{2}$ [see Eqs. (2)-(4) below].
@@
As in Ref. [53], the effective entry and exit rates of particles into and out of the TASEP lanes are parametrized by the parameters $\alpha_{i}$ and $\beta_{i}$ ($i=1,2$), which are positive definite, as follows:
@@
The functions $f(N_{i})$ and $g(N_{i})$ must be monotonically increasing and decreasing in $N_{i}$, respectively, since an increasing reservoir population promotes particle entry into the associated TASEP lane but inhibits their exit. Various forms of such rate functions satisfying monotonicity have been proposed in the literature [57, 42, 48, 49, 58, 50, 51, 52]. For reasons of simplicity, we use the forms in Ref. [53]
@@
To reduce the number of control parameters, we consider the symmetric case $\alpha_{1}=\alpha_{2}\equiv\alpha$ and $\beta_{1}=\beta_{2}\equiv\beta$. These choices are distinct from those in Ref. [53], and allow us to explore a reduced parameter space different from the one considered in Ref. [53]. The model is thus governed by three parameters: $\alpha$, $\beta$, and $\mu$.
@@ III MEAN-FIELD THEORY
In this section, we analyze the model at the mean-field [56] level, following the logic developed in Ref. [53].
@@
Equations (6)-(8) are invariant under the following set of transformations (see Appendix A for details):
@@
In order to set up a continuum description, we introduce a spatial variable $x=j\epsilon$, with lattice spacing $\epsilon=1/L$. In the limit $L\gg 1$, $x$ becomes quasi-continuous, and the density $\rho^{(i)}(x,t)$ is labeled accordingly.
@@
becomes constant. Denoting the constant currents in lanes $T_{1}$ and $T_{2}$ as $J^{(1)}$ and $J^{(2)}$ respectively, the flux balance equation relating these currents to the reservoir occupation numbers $N_{1}$ and $N_{2}$ reads:
@@ IV MEAN-FIELD PHASE DIAGRAMS OF THE MODEL FOR $0<\mu\leq 1$
Before we analyze the phase diagrams of the present model, it is useful to review of the phase diagram of an open TASEP [59, 56, 60]. In an open TASEP, depending on the entry or exit rates ($\alpha_{T}$ and $\beta_{T}$), one finds LD, HD, and MC phases with (constant) steady state densities $\rho_{\text{LD}}=\alpha_{T}$, $\rho_{\text{HD}}=1-\beta_{T}$, and $\rho_{\text{MC}}=1/2$, respectively, in the bulk. Conditions for obtaining these phases are: $\alpha_{T}<\beta_{T}$ and $\alpha_{T}<1/2$ (LD phase), $\beta_{T}<\alpha_{T}$ and $\beta_{T}<1/2$ (HD phase), and $\alpha_{T},\beta_{T}\geq 1/2$ (MC phase). In addition to these, a delocalized domain wall (DDW) is also formed for $\alpha_{T}=\beta_{T}<1/2$. This DDW is manifested as an inclined line spanning the entire TASEP lane and is a consequence of the particle nonconserving dynamics in an open TASEP. Now, transitions between different phases can be identified by equating the particle currents in the associated phases. Thus, the boundaries between either LD or HD and MC phases come out to be straight lines parallel to the $\beta_{T}$ and $\alpha_{T}$ axes respectively, wherein the LD and HD phases are separated by a coexistence line (CL) which starts at (0,0) and terminates at (1/2,1/2) in the $\alpha_{T}$-$\beta_{T}$ plane. Considering the bulk density, $\rho$, as the order parameter, one finds the transition from LD to HD phase as discontinuous change in $\rho$, thereby implying it a discontinuous phase transition. Contrary to this, the transition from either LD or HD to MC phase is seen to be a continuous change in $\rho$, implying it a continuous phase transition.
@@
We now illustrate the phase diagrams of our model in the plane of control parameters $\alpha$ and $\beta$ for a set of representative values of filling factor $\mu$ (cf. Fig. 2 and 10). We use mean-field theory to obtain the phase diagrams and density profiles, which are corroborated by an extensive Monte Carlo simulation (MCS) studies with random sequential updates. We note that more ordered update schemes, such as parallel (synchronous) updates, have also been widely used in the context of traffic and transport models; see, for instance, Refs. [61, 62], which may lead to a simpler, deterministic bulk dynamics. We restrict ourselves here to using random sequential updates, which are well-suited to capturing the steady states. MFT and MCS results agree well with each other. In our model, depending on the value of $\mu$, two or four distinct phases appear simultaneously. These are LD-LD, HD-HD, MC-MC and DW-DW phases. Interestingly, the domain wall solutions, which form the DW-DW phase in our model, always emerge in pair for a particular set of values of $\alpha,\,\beta$, and $\mu$ are actually moving density shocks, not pinned to specific locations in the TASEP lanes, having envelops spanning a finite region across the lanes. These are the delocalized domain walls (DDW) mentioned above. For small values of $\mu$, such as $\mu=3/8$ [Fig. 2(a)], only the LD-LD and DW-DW phases are present, separated by a curved phase boundary. This structure persists for all $0<\mu<1/2$, with the LD-LD region gradually shrinking as $\mu$ increases. In this regime, the HD-HD and MC-MC phases are absent. As $\mu$ increases beyond $1/2$, the MC-MC and HD-HD phases start to emerge, for instance see the phase diagram at $\mu=5/8$ [Fig. 2(b)]. The LD-LD and HD-HD regions now share boundaries with the MC-MC phase, and these boundaries are straight lines parallel to the $\beta$ and $\alpha$ axes, respectively, independent of the value of $\mu$. This topological structure remains unchanged at $\mu=7/8$ [Fig. 2(c)], where the HD-HD region continues to grow while the LD-LD and DW-DW regions contract. At the half-filling point ($\mu=1$) [Fig. 2(d)], the phase diagram simplifies further. All the phase boundaries become straight lines, aligned with the coordinate axes in the $\alpha$-$\beta$ phase diagram. Moreover, the particle-hole symmetry of the model is used to construct the phase diagrams above half-filling ($\mu>1$) from their counterparts below half-filling; see Fig. 10 in Appendix A. From these, it is evident that as $\mu$ increases, the LD-LD region diminishes while the HD-HD and DW-DW regions expand. For $\mu>3/2$, both the LD-LD and MC-MC phases vanish, leaving only the HD-HD and DW-DW phases, a trend that continues up to $\mu=\mu_{\text{max}}=2$. Thus, for $1/2<\mu<3/2$, all four phases — LD-LD, HD-HD, MC-MC, and DW-DW — coexist and meet at a single point known as the four-phase multicritical point.